A stand-by generator is connected to fifteen 40w lamps and a musician's 600w amplifying system, how much energy is used if the generator runs for 6hrs?
1
Expert's answer
2017-09-27T10:43:10-0400
∆E=A=P*t=(15*P_lamp+P_amp )t=(15*40+600)*6*60*60=2.592*10^7 J